At 9AM, a freight train leaves Boston for NYC, traveling at an average rate of 24 miles per hour. At noon, a passenger train sets out on the same route, traveling at an average rate of 60 miles per hour. How far from Boston do the two trains pass each other?
Found 2 solutions by Alan3354, solver91311:
Answer by Alan3354(69443)

At 9AM, a freight train leaves Boston for NYC, traveling at an average rate of 24 miles per hour. At noon, a passenger train sets out on the same route, traveling at an average rate of 60 miles per hour. How far from Boston do the two trains pass each other?
=============
At noon, the freight train is 72 miles away (24*3).
---
The Pass train "gains on" it at 36 mi/hr (60 - 24).
They meet at 72/36 = 2 hours.
==============
Distance from BOS = 2*60 = 5*24 = 120 miles

Answer by solver91311(24713)



The freight train gets a 72-mile head start because it left three hours before the passenger train going 24 miles per hour.

In order for the two trains to pass, the freight train is going to travel some distance in some amount of time going 24 mph, which is to say:



And the passenger train has to go that same distance plus the 72-mile head start going 60 mph, in other words:



Solve the 2X2 system for .

John

My calculator said it, I believe it, that settles it